I am so confused on how your supposed to do this type of substitution problem.
y=-2x+6
3y-x+3=0

Answer by mathsmiles(68)

What you have here are two equations with two variables, X and Y.
You want to substitute so that you end up with only one variable in an equation, which you can then solve.

Lucky for you, they already put one of the equations in a form so that you have Y by itself.
Y = -2X+6
Take this and substitute this for Y in the second equation:
3Y-X+3=0 so ...
3(-2X+6) - X + 3 = 0
Multiplying out the parens:
-6X + 18 - X + 3 = 0
Now combine like terms so ...
(-6X - X) + (18 + 3) = 0
-7X + 21 = 0
Add 7X to both sides of the equation:
21 = 7X
Divide by 7 on both sides:
3 = X

Now we need to figure Y:
Y = -2(3) + 6
Y = -6 + 6
Y = 0

Checking:
3Y - X + 3 = 0
3(0) - 3 + 3 = 0
0-3+3=0 Correct!